n = 81 s2 = 625

H0: σ2 = 500
Ha: σ2 ≠ 500

9. The test statistic equals
a. 100
b. 101.88
c. 101.25
d. 64
10. The p-value is between
a. 0.025 and 0.05
b. 0.05 and 0.1
c. 0.1 and 0.2
d. 0.2 and 0.3
11. What is your conclusion? Use α = 0.05.
a. fail to reject the null hypothesis
b. reject the null hypothesis
Answer #1

9)
Test statistic,
Χ^2 = (n-1)*s^2/σ^2
Χ^2 = (81 - 1)*625/500
Χ^2 = 100

option A

10)
P-value = 0.1291

Option C

11)
Option A